A can do a piece of work in 3 days B can do it in 4 1l2 days.how long will they take to complete the work if they work together?sir pls give me its full steps?
Answer by hamsanash1981@gmail.com(151)

A can do a piece of work in 3 days and B in 4 1/2 ie 9/2 days then,
A's one day work is 1/3 and B's one day work is 1/(9/2) => 2/9
A+B 's one day work = 1/3 + 2/9
= (3+2)/9
= 5/9
A+B can complete the work in 9/5 days or 1 4/5 days